OpenStudy (anonymous):

*I will give medal* The rectangular prism shown has a volume of 140 m3. This prism is enlarged using a scale factor of 2. What is the volume of the enlarged prism?

OpenStudy (anonymous):

help me help you

OpenStudy (anonymous):

I messed up, give me a sec.

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

5*2=10 5*2=10 5.6*2=11.2\[10\times10\times11.2=1120m ^{3}\]

OpenStudy (anonymous):

how do i give you a medal?

OpenStudy (anonymous):

Click "Best Response" located on my comment.

OpenStudy (anonymous):

ok then i think i already gave you one :) teehee